The numbers of workers of a new fast-food restaurant with 7 outlets are 8, 9, 8, 10, 12, 11, and 7. compute the following: Population mean: Population variance population standard deviation


1
Expert's answer
2022-03-17T03:41:38-0400

"\mu=\frac{1}{7}\Sigma_{i=1}^7x_i=\frac{65}{7}=9.2857."

"\sigma^2=\frac{1}{7}\Sigma_{i=1}^7(x_i-\mu)^2=\frac{19.4286}{7}=2.7755."

"\sigma=\sqrt{2.7755}=1.6660" .